Prove (m+n)(p+q) = (mp+np) + (mq+nq)
I have already proved a proposition that states (m+n)p = mp + np, which matches up with part of the right side of the equation, but I'm not sure how to get to that point or how to incorporate mq+nq into it.

Answer by richard1234(7193)

(m+n)q = mq+nq

By the distributive property, (m+n)(p+q) = (m+n)p + (m+n)q = (mp+np) + (mq+nq)